# RealMAN: A Real-Recorded and Annotated Microphone Array Dataset for Dynamic Speech Enhancement and Localization

### v1

@@ 3.4 Data annotation
Direct-path target clean speech. Deep learning-based speech enhancement (dereverberation and denoising) methods require a target clean signal for training. Normally, the direct-path speech signal is used [28, 29]. For simulated datasets, the direct-path speech can be directly generated using the simulated direct-path impulse response. For real-recorded datasets, providing real direct-path speech has never been solved in the field. Instead of providing a target signal that can be used for training, existing real-recorded datasets either provide speech transcriptions or close-talk speech signals [19, 20, 21, 23].
@@
In this paper, we develop a method to estimate the direct-path speech by source speech (replayed speech) and the microphone recordings. The recording process is formulated in the time domain as $x(t)=s_{dp}(t)+s_{rev}(t)+n(t)$, $s_{dp}(t)=h_{dp}(t)*h_{dev}(t)*s(t)=h_{dev}*\left[As(t-\tau)\right]$, where $x(t)$, $s_{dp}(t)$, $s_{rev}(t)$ and $n(t)$ are the microphone recording, direct-path speech, speech reverberation, and noise, respectively. Theoretically, the direct-path speech is the convolution of the played source speech $s(t)$, the impulse response of the playing and recording devices $h_{dev}(t)$, and the direct-path impulse response $h_{dp}(t)$, where the direct-path impulse response $h_{dp}(t)$ can be formulated as a level attenuation $A$ and a time shift $\tau$ of $s(t)$. Note that $A$ and $\tau$ are time-invariant for static source, while time-varying for moving source. The impulse response of devices $h_{dev}$ can be measured in advance. Then, the estimation of ${s}_{dp}(t)$ amounts to the estimation of $A$ and $\tau$ according to the known $x(t)$, $h_{dev}$, and $s(t)$. For details of the algorithm, readers can refer to Appendix C.1.
By informal listening test, the estimated direct-path speech has the same perceptual quality as source speech. In addition, to further evaluate the potential impact of the proposed method, we test the ASR performance with the established ASR model trained on over 10,000 hours of Mandarin dataset WenetSpeech [30] using the ESPNet toolkit. The character error rates (CERs) for the source speech and the estimated direct-path speech are identical, which means the direct-path filtering of source speech do not degrade the intelligibility of speech. Overall, the estimated direct-path speech could be a good target signal for training speech enhancement networks. The intermediate results shown in Appendix C.1 also strongly support our claims.
@@
Sound source location. The annotation of source (loudspeaker) location leverages a fisheye camera and a LED light (placed on top of the loudspeaker). During the speech recording process, the fisheye camera is placed right above the microphone array, and the plane coordinate of the camera and microphone array are aligned. Accordingly, the azimuth angle of source could be calculated by the position of the LED light in the camera image. To locate the LED light in the image, a visual-based LED light is designed and applied on each recorded video frame. To guarantee the accuracy, all the detection results are manually checked. Please refer to Appendix C.3 for the illustration examples and pseudocodes of the LED detection algorithm. According to the frame rate of the fisheye camera, the temporal resolution of the annotation of source location is set to 100 ms.
@@ 3.5 Dataset split and statistics
Speech duration (h) 63.50 7.78 11.64
     - Moving speaker (h) 26.57 3.27 4.75
     - Static speaker (h) 36.93 4.51 6.89
Noise duration (h) 106.32 15.95 22.18
@@
The recorded speech and noise data are split into training, validation, and test sets for the deep learning-based speech enhancement and source localization methods, according to the acoustic characteristics of the recording scenes and speaker identities:
Acoustic characteristics of scenes. Different scenes have different RIRs and noise characteristics. To make sure that the model can be trained under diverse scenes, there are 40 different scenes (of speech and noise) included in the training set. Various types of acoustic scenes are also provided in the validation and test sets (17 and 21, respectively), such that the algorithms can be fully evaluated under various scenarios. There are 3 scenes that only appeared in the test and validation sets, respectively, to further evaluate the generalization capability on the unseen acoustic scenes. Note that, although some scenes may overlap across sets, there is no data sample overlap among them.
Speaker identities. The entire 55 speakers are splitted into 43, 6, and 6 for the training, validation and test sets, respectively. Following the general speech corpus split, no speaker appears across sets.
Speech and noise matching. During recording, speech and noise are normally recorded in quiet midnight and noisy daytime, respectively. To make the noisy speech as real as possible, it is better to mix speech and noise from the same scene. This principle is followed for validation and test sets. The number of scenes with both speech and noise recordings is 10 out of 17 for validation scenes and 11 out of 21 for test scenes. The same type of indoor environments, such as living rooms or office rooms, have similar noise characteristics, hence the noises are not recorded for every scene. For these cases, we mix the speech of one scene with noise from a similar environment. Specifically, speech of all ClassRooms are mixed with ClassRoom1 noise; speech of all OfficeRooms and Library are mixed with the noises of OfficeRoom1/3; speech of all LivingRooms are mixed with noises of LivingRoom1 and Laundry. As for training, some preliminary experiments show that such match on speech and noise scenes is not required. And a random scene match between speech and noise is more suitable for network training, which is possibly because of the promotion on data diversity.
The statistics of the dataset are shown in Table 3. The total 83.9 hours of the recorded speech are divided into 63.5, 7.8 and 11.6 hours for training, validation and test, respectively. And 144.5 hours of noise data are divided into 106.3, 16.0 and 22.2 hours for training, validation and test, respectively.
@@ 4 Baseline Experiments
In this section, we benchmark the proposed dataset for speech enhancement and source localization. As presented in Section 3.5, the training set is generated by randomly mixing speeches and noises, with a signal-to-noise ratio (SNR) uniformly distributed in [0, 15] dB. By contrast, the validation and test sets are generated by mixing speeches and noises from matched scenes, and the signal level of mixed speech and noise are kept unchanged as recorded to maintain their natural loudness.
@@ 4.1 Baseline methods and evaluation metrics
Speech enhancement. One popular time-domain network, i.e. FaSNet-TAC [10], and one recently-proposed frequency-domain network, i.e. SpatialNet [28], are used for benchmarking the speech enhancement performance of the proposed dataset. The negative of scale-invariant signal-to-distortion ratio (SI-SDR) [31] is used as the loss function for training the two baseline networks. For FaSNet-TAC, the best configuration reported in its original paper is used. For SpatialNet, to reduce the computational complexity, a tiny version is used, where the hidden size of the SpatialNet-small version reported in the paper [28] is further reduced from 96 to 48.
SI-SDR, WB-PESQ [32], and MOS-SIG, MOS-BAK, MOS-OVR from DNSMOS [33] measures the speech enhancement performances. The ASR performances are evaluated by an established ASR model trained by over 10,000 hours of Mandarin dataset, WenetSpeech [30] using ESPNet toolkit.
Sound source localization. Azimuth angle localization is performed. We adopt a convolutional recurrent neural network (CRNN) as one baseline system for sound source localization. The baseline CRNN comprises a 10-layer CNN and a 1-layer gated recurrent unit. The kernel size of convolutional layers are all $3\times 3$, each convolutional layer is followed by an instance normalization and a rectified linear unit activation function. Max pooling is applied to compress the frequency and time dimensions after every two convolutional layers. This baseline CRNN is very similar to the CRNN network used in many sound source localization methods [34, 35, 36]. The spatial spectrum, with candidate locations of every $1^{\circ}$ azimuth angle, is used as the learning target [8, 37, 38, 39]. A linear classifier with sigmoid activation is used to predict the spatial spectrum. A recently proposed sound source localization method, i.e. IPDnet [12], is also used a baseline system. The hidden size of the original IPDnet is reduced from 256 to 128. Candidate locations are also set as every $1^{\circ}$ azimuth angle.
The localization results are evaluated with i) the Mean Absolute Error (MAE) and ii) Localization Accuracy (ACC) ($N^{\circ}$), namely the ratio of frames with the estimation error of azimuth less than $N^{\circ}$.
@@
unprocessed - 1.15 -9.8 2.00 1.72 1.51 19.9 1.11 -9.1 1.79 1.54 1.36 23.8
FaSNet-TAC [10] sim sim 1.38 -3.4 2.67 3.19 2.22 27.1 1.33 -2.5 2.60 3.12 2.14 29.7
sim real 1.49 -1.7 2.83 3.23 2.35 22.4 1.42 -1.5 2.78 3.20 2.29 25.7
real sim 1.47 0.8 2.67 3.09 2.18 23.7 1.40 0.5 2.58 3.05 2.10 28.2
real real 1.51 1.3 2.80 3.34 2.35 22.4 1.43 1.1 2.73 3.28 2.27 26.3
SpatialNet [28] sim sim 1.40 -8.4 3.09 2.62 2.28 19.2 1.33 -7.9 3.06 2.53 2.23 23.2
sim real 1.45 -2.6 2.58 2.35 1.95 23.0 1.38 -2.6 2.54 2.25 1.89 26.5
real sim 1.96 3.8 3.09 3.06 2.45 17.3 1.80 3.0 3.00 2.99 2.36 21.2
real real 2.10 6.1 3.05 3.51 2.62 16.0 1.90 3.8 2.96 3.45 2.52 21.5
@@ 4.2 Benchmark experiments
This section benchmarks the speech enhancement and source localization performance of the proposed dataset, using a 9-channel sub-array (the smallest circle and the center microphone, and the center microphone is used as the reference microphone when necessary). In addition, we also evaluate the effect of the simulation-to-real mismatch on speech enhancement and source localization tasks. Equal amounts of multichannel speech are simulated according to our real-recorded dataset. Specifically, one counterpart utterance is simulated for each real-recorded utterance using the same source speech, room size, T60, and source position/trajectory as the real-recorded utterance. Note that source speech, room size, and T60 of each real-recorded utterance are available in our dataset. The source position/trajectory can be obtained with the calibration parameters of the fisheye camera. However, the elevation angle and source distance are not very accurate, which we think is not problematic for generating the simulated speech data, but the elevation angle and source distance will not be publicly released. Multichannel noise is simulated with the diffuse noise generator [5], using single-channel of white, babble, and factory noise.
The simulated speech/noise can also be combined with real-recorded noise/speech. For each setting, the signal-to-noise (SNR) for mixing speech and noise is uniformly sampled in [0, 15] dB.
@@
Speech Enhancement. The results of speech enhancement are shown in Table 4. Overall, compared to other settings, training with real speech and real noise achieves the best speech enhancement performance on the real-recorded test set, for both the baseline networks. As for intrusive metrics, i.e. WB-PESQ and SI-SDR, the target clean speech provided in this dataset are used as the reference signals, which may leads to some measurement bias for other settings, therefore these metrics are only presented for reference. The non-intrusive DNS-MOS scores can better reflect the speech quality. It can be seen that, for FaSNet-TAC, training with simulated speech and real noise achieves almost the same DNS-MOS performance as the setting of real speech plus real noise, which indicates that speech simulation does not have the simu-to-real problem for speech enhancement. However, this is not the case for SpatialNet, for which training with simulated speech and real noise achieves the worst performance. Some validation experiments have been conducted to figure out the reasons for this phenomenon, and the most possible reason is: there are some slight mismatches between the real and designing microphone positions, which lead to severe overfitting when training with simulated speech and real noise due to some unclear factors. These reflect that different networks exploit different information for speech enhancement, and perform differently under certain conditions as well.
@@
Training with real speech and real noise consistently outperforms training with simulated noise. The simulated noise lacks diversity in terms of both spectral pattern and spatial correlation. The spatial correlation of real noise could largely vary from the one of theoretical diffuse noise field. Moreover, the spatial correlation of real noise is also highly time-varying. Please see Appendix D.2 for some real examples. The mismatch between diffuse noise and real noise leads to the performance degradation when training with simulated noise. In addition, due to high complexity and non-stationarity of the spatial correlation of real noise, it is complicated to develop new techniques to simulate real noise. Therefore, we suggest to use real-recorded noise for training speech enhancement networks.
@@
Sound source localization. The results of the sound source localization are presented in Table 6. It can be observed that both real speech and real noise are beneficial for improving the sound source localization performance. Different from the speech enhancement task using simulated speech may not be problematic, the mismatch between real recordings and simulated RIRs causes the performance degradation of sound source localization, which is consistent with the findings in [40]. The mismatch between simulated and real noise has a great influence as well.
@@
sim sim 71.9 10.2 68.8 9.6
sim real 76.7 9.9 70.3 11.1
real sim 82.1 8.1 75.9 8.2
real real 88.4 4.6 83.9 4.3
@@
Fixed-Array[12] 86.1 3.6 88.9 2.7
Variable-Array[12] 86.1 3.5 80.4 3.6

### v2

@@ 3.4 Data annotation
Direct-path target clean speech. Deep learning-based speech enhancement methods require a target clean signal for training. Normally, the direct-path speech signal is used [34, 35]. For real-recorded datasets, before, providing direct-path speech has never been solved in the field.
@@
In this paper, we develop a method to estimate the direct-path speech based on the source speech (replayed speech) and microphone recordings. The recording process is formulated in the time domain as $x(t)=s_{dp}(t)+s_{rev}(t)+n(t)$, $s_{dp}(t)=h_{dp}(t)*h_{dev}(t)*s(t)=h_{dev}*\left[As(t-\tau)\right]$, where $x(t)$, $s_{dp}(t)$, $s_{rev}(t)$ and $n(t)$ are the microphone recording, direct-path speech, speech reverberation, and noise, respectively. Theoretically, the direct-path speech is the convolution of the played source speech $s(t)$, the impulse response of the playing and recording devices $h_{dev}(t)$, and the direct-path impulse response $h_{dp}(t)$, where the direct-path impulse response $h_{dp}(t)$ can be formulated as a level attenuation $A$ and a time shift $\tau$ of $s(t)$. Note that $A$ and $\tau$ are time-invariant for static source, while time-varying for moving source.
The impulse response of devices (microphone and loudspeaker combined) $h_{dev}$ is considered to be constant (independent to the direction and orientation of the loudspeaker relative to the microphone) and it is measured in advance for one configuration where the loudspeaker faces toward the microphone with a source-to-microphone distance of 1 m, please refer to Appendix B.3 for more details. This simplified measurement is accurate for the omni-directional microphone (except for a constant time delay and level factor), but inaccurate for the frontal-directional loudspeaker as its impulse response is orientation-dependent. However, in our setting, it is difficult to annotate the loudspeaker orientation and thus to take the loudspeaker directivity into account, which is left for future research.
Then, the estimation of ${s}_{dp}(t)$ amounts to the estimation of $A$ and $\tau$ according to the known $x(t)$, $h_{dev}$, and $s(t)$. Note that, any constant time delay and level factor, such as the one caused by the measurement of $h_{dev}$ or by the loudspeaker orientation, will be absorbed into the estimated $\tau$ and $A$, respectively. For details of the estimation algorithm, please refer to Appendix C.1. The direct-path target speech can be estimated in the same way for all microphone channels. To keep a small data size, only the direct-path target speech for microphone 0 is included in the released dataset. In the future, we can provide more if there is a high requirement.
Due to the lack of ground truth, it is not straightforward to evaluate the estimation accuracy. We think the credibility of the estimated direct-path speech can be well testified based on the following criteria. i) The estimated direct-path speech should have the same speech quality as source speech to provide an ideal upper-bound for speech enhancement. We have conducted an informal listening test by eight listeners to test whether there is an audible difference between the perceptual quality of the estimated speech and source speech, which shows that there is no audible difference with a 91% confidence. We have also tested the ASR performance, with an established ASR model trained on over 10,000 hours of Mandarin dataset WenetSpeech [36], where the results are identical for the estimated direct-path speech and source speech. ii) As shown in Appendix C.1, the estimation of $A$ and $\tau$ is based on the cross-correlation method. The good properties of intermediate results, such as the sharp correlation peak and the smooth estimation curve for moving source, also give us strong confidence on the estimation accuracy. iii) After all, as long as speech enhancement networks can be successfully trained with the estimated direct-path speech as the target (will be shown in the experiment section), the estimated direct-path speech can be deemed to be sufficiently accurate.
Sound source location. The annotation of source/loudspeaker location leverages a fisheye camera and an LED light (placed on top of the loudspeaker). During the speech recording process, the fisheye camera is placed right above the microphone array, and the plane coordinates of the camera and microphone array are aligned. The height of loudspeaker center (varying from 1.3 m to 1.6 m) is manually logged in the recording process. Given the pixel position of the LED light in the camera image, the source location can be calculated in three steps. i) Based on camera calibration, the azimuth angle (relative to both the array and camera) and elevation angle (relative to only the camera) of source can be calculated with the pixel position. ii) With the height and elevation angle (relative to the camera) of source, the horizontal distance between the source and camera/array can be calculated. iii) The elevation angle (relative to the array) of source can be calculated using the height of microphone array, the height and the horizontal distance of source. Note that, the fixed 12 cm height difference between the LED light and loudspeaker center has been taken into account.
@@
To LED light is detected for each video frame with a vision-based detection algorithm. To guarantee the accuracy, all the detection results are manually double-checked. Please refer to Appendix C.3 for examples and pseudocodes of LED detection. According to the frame rate of the fisheye camera, the temporal resolution of source location annotation is set to 100 ms.
@@ 3.5 Dataset split and statistics
Speech duration (hour) 64.0 8.1 11.6
     - Moving speaker (hour) 27.1 3.5 4.8
     - Static speaker (hour) 36.9 4.6 6.8
Noise duration (hour) 106.3 16.0 22.2
@@
The recorded speech and noise data are split into training, validation, and test sets for learning-based speech enhancement and source localization, according to the acoustic characteristics of the recording scenes and speaker identities:
Acoustic characteristics of scenes. Different scenes have different RIRs and noise characteristics. To make sure that the model can be trained under diverse scenes, there are 40 different scenes (of speech and noise) included in the training set. Various types of acoustic scenes are also provided in the validation and test sets (17 and 21, respectively), such that the algorithms can be fully evaluated under various scenarios. There are 3 scenes that only appeared in the test and validation sets, respectively, to further evaluate the generalization capability on the unseen acoustic scenes. We think the scene diversity of training and test sets are both critical for the full evaluation of general-purpose speech enhancement and source localization methods, so we have to make some scene overlaps across sets, but note that there is no data sample overlap among them.
Speaker identities. Following the general speech corpus split, the entire 55 speakers are split into 43, 6, and 6 for the training, validation, and test sets, respectively, without speaker overlap across sets. Please refer to Appendix B.2 for detailed speech duration statistics across speakers.
Speech and noise matching. During recording, speech and noise are normally recorded in quiet midnight and noisy daytime, respectively. To make the noisy speech as real as possible, it is better to mix speech and noise from the same scene. This principle is followed for the validation and test sets. The number of scenes with both speech and noise recordings is 10 out of 17 for validation scenes and 11 out of 21 for test scenes. The same type of indoor environments, such as living rooms or office rooms, have similar noise characteristics, hence noise is not recorded for every scene. For these cases, we mix the speech of one scene with noise from a similar environment. Specifically, speech of all ClassRooms are mixed with ClassRoom1 noise; speech of all OfficeRooms and Library are mixed with OfficeRoom1/3 noise; speech of all LivingRooms are mixed with LivingRoom1 and Laundry noise. As for training, some preliminary experiments show that such match on speech and noise scenes is not required. Instead, a random scene match between speech and noise is more suitable for network training, possibly because of the promotion of data diversity.
The statistics of the dataset are shown in Table 3. The total 83.7 hours of the recorded speech are divided into 64.0, 8.1, and 11.6 hours for training, validation, and test, respectively. The total 144.5 hours of noise data are divided into 106.3, 16.0 and 22.2 hours for training, validation and test, respectively.
@@ 4 Baseline Experiments
In this section, we benchmark the proposed dataset for speech enhancement and source localization. As presented in Section 3.5, the validation and test sets are generated by mixing speeches and noises from matched scenes, and the signal level of mixed speech and noise are kept unchanged as recorded to maintain their natural loudness. The speech-noise mixed validation and test sets have an average SNR about 0.0 dB and -0.8 dB, respectively. Please refer to Appendix B.5 for the SNR statistics of validation and test sets. The training set is generated by randomly mixing speeches and noises, with a speech-recording-to-noise-recording ratio (SNR) uniformly distributed in [-10, 15] dB. In Appendix D.3, we also provide the experimental results on the recorded speech without adding noise.
In this paper, we only perform single-speaker speech enhancement (denoising and dereverberation) and source localization. Nevertheless, we think the proposed dataset can also be used for the tasks of multi-speaker separation and localization. Multi-speaker signals can be generated by simply mixing the single-speaker signals recorded in the same scene, which will be highly consistent to the simultaneous recording of multiple speakers. One unreal factor is that the background noise in speech recordings will also be mixed, which we think is not problematic, as the SPL of background noise is low compared to the SPL of speech, i.e. 57 dB (36 dBA) versus 68 dB (61 dBA).
@@ 4.1 Baseline methods and evaluation metrics
Speech enhancement. One popular time-domain network, i.e. FaSNet-TAC [13], and one recently-proposed frequency-domain network, i.e. SpatialNet [34], are used for benchmarking the speech enhancement performance. The negative of scale-invariant signal-to-distortion ratio (SI-SDR) [37] is used as the loss function for training the two networks. For FaSNet-TAC, the best configuration reported in its original paper is used. For SpatialNet, to reduce the computational complexity, a tiny version is used, where the hidden size of the SpatialNet-small version reported in the paper [34] is further reduced from 96 to 48. SI-SDR (in dB), WB-PESQ [38], and MOS-SIG, MOS-BAK, MOS-OVR from DNSMOS [39] are used for measuring the performance of speech enhancement. The ASR performance are evaluated by the WenetSpeech [36] ASR model trained by over 10,000 hours of Mandarin dataset, implemented in the ESPNet toolkit. Character error rate (CER, in percentage) is taken as the ASR metric.
Sound source localization. Azimuth angle localization is performed. We adopt a convolutional recurrent neural network (CRNN) as one baseline system for sound source localization. The baseline CRNN comprises a 10-layer CNN and a 1-layer gated recurrent unit. The kernel size of convolutional layers are all $3\times 3$, each convolutional layer is followed by an instance normalization and a rectified linear unit activation function. Max pooling is applied to compress the frequency and time dimensions after every two convolutional layers. This baseline CRNN is very similar to the CRNN network used in many sound source localization methods [40, 41, 42]. The spatial spectrum, with candidate locations of every $1^{\circ}$ azimuth angle, is used as the learning target [10, 43, 44, 45]. A linear classifier with sigmoid activation is used to predict the spatial spectrum. A recently proposed sound source localization method, i.e. IPDnet [15], is also used a baseline system. The hidden size of the original IPDnet is reduced from 256 to 128. Candidate locations are also set as every $1^{\circ}$ azimuth angle. IPD templates are computed with the theoretical time delays from candidate locations to microphones, following [15]. The training target of ground truth direct-path IPDs are computed with the annotated azimuth angles. The localization results are evaluated with i) the Mean Absolute Error (MAE) and ii) Localization Accuracy (ACC) ($N^{\circ}$), namely the ratio of frames with the estimation error of azimuth less than $N^{\circ}$. Please refer to Appendix D.1 for more detailed experimental configurations.
@@
unprocessed - 1.14 -9.9 2.01 1.73 1.52 20.1 1.10 -9.0 1.80 1.54 1.37 23.7
FaSNet-TAC [13] sim sim 1.39 -2.4 2.71 3.00 2.18 25.4 1.34 -2.5 2.62 2.92 2.10 28.2
sim real 1.47 -1.0 2.83 3.24 2.36 20.5 1.41 -0.8 2.75 3.12 2.27 23.4
real sim 1.42 0.4 2.63 3.21 2.19 24.0 1.36 0.4 2.51 3.15 2.08 28.2
real real 1.46 2.1 2.79 3.30 2.34 21.9 1.39 1.2 2.73 3.20 2.26 25.6
SpatialNet [34] sim sim 1.37 -5.2 3.24 2.85 2.46 17.9 1.37 -4.8 3.20 2.75 2.40 21.0
sim real 1.13 -9.7 1.59 1.67 1.32 44.2 1.11 -9.2 1.51 1.57 1.26 49.0
real sim 1.96 4.7 3.16 3.13 2.52 17.3 1.80 2.7 3.08 3.05 2.42 21.2
real real 2.16 7.3 3.23 3.43 2.71 14.5 1.97 4.3 3.15 3.40 2.63 18.6
@@ 4.2 Benchmark experiments
Benchmark experiments use a 9-channel sub-array (the smallest circle and the center microphone, and the center microphone is used as the reference microphone when necessary). In addition, we also evaluate the effect of the simulation-to-real mismatch on speech enhancement and source localization tasks. Equal amounts of multichannel speech are simulated according to our real-recorded dataset, using the gpuRIR toolkit [46]. Specifically, one counterpart utterance is simulated for each real-recorded utterance using the same source speech, room size, T60, and source position/trajectory as the real-recorded utterance. The directivity of microphone and source are always set to be omni-directional during simulation. Multi-channel noise is simulated with the diffuse noise generator [5], taking white (code generated), babble and factory noise (from noisex-92 [47]) as the single-channel source noise. The simulated speech/noise can also be combined with real-recorded noise/speech. For each setting, the SNR for mixing speech and noise is uniformly sampled in [-10, 15] dB.
@@
Speech Enhancement. The results of speech enhancement are shown in Table 4. Overall, compared to other settings, training with real speech and real noise achieves the best speech enhancement performance on the real-recorded test set, for both the baseline networks. As for intrusive metrics, i.e. WB-PESQ and SI-SDR, the target clean speech provided in this dataset are used as the reference signals, which may leads to some measurement bias for other settings, therefore these metrics are only presented for reference. The non-intrusive DNS-MOS scores can better reflect the speech quality. It can be seen that, for FaSNet-TAC, training with simulated speech and real noise achieves comparable DNS-MOS performance as the setting of real speech plus real noise, which indicates that speech simulation does not have the simulation-to-real problem for FaSNet-TAC. However, this is not the case for SpatialNet, for which training with simulated speech and real noise achieves the worst performance. Some validation experiments had been conducted to figure out the reasons for this phenomenon, which showed that there might be a slight mismatch between the real and ideal microphone positions. We have designed a new method for resolving the problem of microphone position mismatch, namely disturbing the ideal microphone positions in training, which is shown to be very effective on simulated test data, but only slightly improve the performance on our real test data. This indicates that there are still some unclear mismatches between the simulated and real data. As its name indicates, SpatialNet [34] mainly learns the RIR-related spatial information, which is possibly more sensitive to those unclear mismatches. Exploring and resolving those mismatches would be an interesting topic for future research.
@@
Training with real speech and real noise consistently outperforms training with simulated noise. The simulated noise lacks diversity in terms of both spectral pattern and spatial correlation. The spatial correlation of real ambient noise could largely deviates from the one of theoretical diffuse noise field. Moreover, the spatial correlation of real noise is also highly time-varying. Please see Appendix D.2 for the detailed analysis of spatial correlation of real noise. The simulation-to-real mismatch of noise leads to the performance degradation. In addition, due to the high complexity and non-stationarity of the spatial correlation of real noise, it is complicated to develop new techniques for simulating real noise. Therefore, we suggest to use real-recorded multi-channel noise for training speech enhancement networks.
@@
Sound source localization. The results of the sound source localization are presented in Table 6. The mismatch between simulated RIRs and real recordings causes the performance degradation of sound source localization, which is consistent with the findings in [12, 48, 49]. The simulation-to-real mismatch of noise causes the performance degradation of sound source localization as well.
@@
sim sim 60.4 7.8 58.8 9.6
sim real 52.4 22.9 48.4 21.4
real sim 80.0 5.7 75.8 8.4
real real 89.2 2.2 86.7 3.1
@@
Fixed-Array 87.0 3.0 85.4 3.1
Variable-Array 87.1 2.4 78.8 3.7